OpenStudy (anonymous):

Which of these are the intercepts of y = 2x − 6?

OpenStudy (anonymous):

Set x=0, solve y; Set y=0, solve x. Then u have (0,-6),(3,0)

OpenStudy (bradely):

y=2x-6 2x-y=6 Divide each side by 6 (2x/6)+(y/-6)=6/6 (x/3)+(y/-6)=1 It is in the form x/a+y/b =1 x-intercept 3 y-intercept -6 Source: http://www.mathskey.com/question2answer/